Two cards are drawn from a well-shuffled deck of 52 cards without replacement. Determine the probability of each event.
a) The second card is spade.

Answer by solver91311(24713)



Sample Space                 Probability
   ♠     ♠      1/4 *  4/17 =   1/17
   ♠    ♣♥♦     1/4 * 13/17 =  13/68
  ♣♥♦   ♣♥♦     3/4 * 38/51 =  19/34
  ♣♥♦    ♠      3/4 * 13/51 =  13/68

P(♠ 2nd) = 1/17 + 13/68 = 1/4



John

My calculator said it, I believe it, that settles it


I > Ø